Uncle Akeju is 65 years of age and has a life expectancy of 12 years. He wishes to invest N20,000 in an annuity that will make a level payment at the end of each year until his death. If the Interest rate is 8 percent, what income can uncle Akeju expect to receive each year


1
Expert's answer
2021-11-14T17:34:03-0500

Using the present value of an annuity formula, we have:

PV=C"\times" "[\frac{1}{R}-\frac{1}{r\times(1+r)^{t}}]"


"\$20,000=C\times[\frac{1}{0.08}-\frac{1}{0.08\times(1+0.08)^{12}}]"


"C" "=\$2,653.90"